OpenStudy (anonymous):

c2 – 4c = 0 , what are the solutions of the equation?

OpenStudy (anonymous):

c(c-4)=0 c=0 c=4

OpenStudy (campbell_st):

factorise the quadratic \[c^2 - 4c = 0.....so.... c(c - 4) = 0\] then solve

OpenStudy (campbell_st):

correct.